ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

9. Ordinary Differential Equations and Dynamical Systems

1991 MS Classification: Okoya, Samuel S., Obafemi Awolowo University, Nigeria A criterion for solutions of a model quasilinear parabolic equation The purpose of the paper is to investigate the existence (and uniqueness) of global smooth solution of a quasilinear parabolic equation which is motivated by a model problem arising from biophasic, nonlinearly permeable viscoelastic material. The method of upper and lower solutions is used to obtained a necessary condition for global existence.